At Reliable's Trucking,Inc., annual sales are predicted to increase over the next 3 years at a rate of 6% per year. Sales equate to "fleet miles." If Reliable's fleet miles are predicted to reach 4.4 million in 3 years, what is the number of fleet miles today? Round to the nearest tenth of a million.
a) 2.45 million
b)1.5 million
c)6.724 million
d)3.694 million

Answer by Theo(13342)

the general equation is:

f = p * (1+r)^n

f is the future amount.
p is the present amount.
r is the interest rate per payment period.
n is the number of payment periods.

r is equal to 6% per year.
the payment period is in year.
n is equal to 3 years.
f is equal to 4.4 million fleet miles.
p is what you are trying to solve for.

the formula becomes:

4.4 = p * (1.06)^3

divide both sides of this equation by (1.06)^3 to get:

4.4 / (1.06)^3) = p

this becomes:

4.4 / 1.191016 = p which becomes:

p = 3.694324845 million fleet miles.

that would correspond to 3.694 million fleet miles rounded to 3 decimal places which is selection d.